OpenStudy (jenniferjuice):

Trevor is currently 22 years old. He is half of what his dad's age was six years ago. How old is Trevor's Dad?

OpenStudy (jenniferjuice):

54 40 50 44

OpenStudy (whpalmer4):

T (Trevor) = 22. D (Dad) T = 1/2 (D-6) (He is 1/2 what Dad was 6 years ago, when Dad was 6 years younger) 22=1/2 (D-6)

OpenStudy (jenniferjuice):

what does d represent?

OpenStudy (jenniferjuice):

wait nevermind

OpenStudy (jenniferjuice):

but would i distribute 1/2?

OpenStudy (whpalmer4):

\[22 = \frac{D-6}{2}\]

OpenStudy (jenniferjuice):